The density of mercury is 13.5 g/mL and the density of water is 1.00 g/mL. If a mercury barometer reads 787 mmHg, what is the barometric pressure in centimeters of water (cmH2O)?
Answer #1

Barometric pressure = d*g*h
where:
d is density
g is acceleration due to gravity
h is height of liquid

SO,
Since Barometric pressure is same for Hg and water:
d(Hg)*h(Hg) = d(H2O)*h(H2O)
13.5 g/mL * 787 mmHg = 1.00 g/mL * h(H2O)
h(H2O) = 10624 mm H2O
h(H2O) = 1062 cm H2O

Answer: 1.06*10^3 cm